OpenStudy (anonymous):

A small manufacturing company makes and sells x machines each month. The monthly cost C, in dollars, of making x machines is given by C (x) = 2600 + 0.4x^2 . The monthly income I, in dollars, obtained by selling x machines is given by I(x)=150x−0.6x^2 . (a) Show that the companyís monthly profit can be calculated using the quadratic function P(x)=−x^2 +150x−2600.

OpenStudy (anonymous):

@FoolForMath

OpenStudy (anonymous):

Profit = Income - Costs, or P(x) = I(x) - C(x)

OpenStudy (anonymous):

I'm guessing there's a typos somewhere b/c the numbers don't work out

OpenStudy (anonymous):

mmmm there's no typo... and not entirely sure how to apply this question to that formula.

OpenStudy (anonymous):

@peachpi

OpenStudy (anonymous):

P(x) = I(x) - C(x) = (150x−0.6x²) - (2600 + 0.4x^2)

OpenStudy (anonymous):

oh and your numbers are right

OpenStudy (anonymous):

dumb mental error on my part

OpenStudy (anonymous):

thanks for all your help

OpenStudy (anonymous):

profit=selling price-cost price so, selling price is 150x-0.6x^2 cost price is 2600 + 0.4x^2